Blood glucose levels for those obese teenagers have a mean of 120. A researcher thinks that a diet high in raw corn starch will have a positive or negative effect on blood glucose levels. A sample of 25 students who have tried the raw corn starch diet has a mean glucose level of 135 with a standard deviation of 38. Test the hypothesis using the level of significance of 0.10 that the raw corn starch diet had an effect.


1
Expert's answer
2022-05-08T13:45:40-0400

1. The following null and alternative hypotheses need to be tested:

"H_0:\mu=120"

"H_a:\mu\not=120"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.


2. Based on the information provided, the significance level is "\alpha = 0.10," "df=n-1=24" degrees of freedom, and the critical value for a two-tailed test is "t_c =1.710882."

The rejection region for this two-tailed test is "R = \{t: |t|>1.710882\}"


3.The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{135-120}{38/\sqrt{25}}=1.973684"



4. Since it is observed that "|t| =1.973684>1.710882= t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=24" degrees of freedom, "t=1.973684" is "p =0.056589," and since "p= 0.056589<0.10=\alpha," it is concluded that the null hypothesis is rejected.


5. Therefore, there is enough evidence to claim that the population mean "\mu" is different than 120, at the "\alpha = 0.10" significance level.